OpenStudy (anonymous):

Please help, will medal and fan \[\frac{ 1 }{ b+1 }+\frac{ 1 }{ b-1 }=\frac{ 2 }{ b ^{2} -1 }\]

OpenStudy (anonymous):

Solve for b

Nnesha (nnesha):

first find common denominator left side

OpenStudy (anonymous):

(b+1)(b-1) ?

OpenStudy (anonymous):

\[\frac{ (b-1)+(b+1) }{ (b-1)(b+1) }=\frac{ 2 }{ b ^{2} -1 }\]

Nnesha (nnesha):

tag meh :P

OpenStudy (anonymous):

is that right? @Nnesha

Nnesha (nnesha):

thanks for tagging i'm not getting not notification :)

Nnesha (nnesha):

yes that is right

Nnesha (nnesha):

now factor denominator of right side

Nnesha (nnesha):

remember difference of square we did yesterday

OpenStudy (idku):

\[\frac{1}{b+1}+\frac{1}{b-1}=\frac{b+1}{(b+1)(b-1)}+\frac{b-1}{(b+1)(b-1)}=\frac{b}{b^2-1}\]

OpenStudy (idku):

\[\frac{b}{b^2-1}=\frac{2}{b^2-1}\]

OpenStudy (anonymous):

wait, @Nnesha did I do it wrong? look at idku's

Nnesha (nnesha):

 this is right now combine like terms at the numerator